A solution of 500-ml was 84% Alcohol. How many gallons of the solution would need to be evaporated to be only 80% alcohol
Answer by stanbon(75887)

A solution of 500-ml was 84% Alcohol. How many gallons of the solution would need to be evaporated to be only 80% alcohol
------
Equation:
alcohol - alcohol = alcohol
0.84*500 - x = 0.80(500-x)
84*500 - 100 x = 80*500 - 80x
4*500 = 20x
x = 100 ml (amt. to evaporate)
=================
Cheers,
Stan H.